### v2

# On the Convergence Rate of Linear $\text{\sf Datalog}^{\circ}$ over Stable Semirings

Abstract
$\text{\sf Datalog}^{\circ}$ is an extension of Datalog, where instead of a program being a collection of union of conjunctive queries over the standard Boolean semiring, a program may now be a collection of sum-product queries over an arbitrary commutative partially ordered pre-semiring. $\text{\sf Datalog}^{\circ}$ is more powerful than Datalog in that its additional algebraic structure alows for supporting recursion with aggregation. At the same time, $\text{\sf Datalog}^{\circ}$ retains the syntactic and semantic simplicity of Datalog: $\text{\sf Datalog}^{\circ}$ has declarative least fixpoint semantics. The least fixpoint can be found via the naïve evaluation algorithm that repeatedly applies the immediate consequence operator until no further change is possible.
It was shown in [12] that, when the underlying semiring is $p$-stable, then the naïve evaluation of any $\text{\sf Datalog}^{\circ}$ program over the semiring converges in a finite number of steps. However, the upper bounds on the rate of convergence were exponential in the number $n$ of ground IDB atoms.
This paper establishes polynomial upper bounds on the convergence rate of the naïve algorithm on linear $\text{\sf Datalog}^{\circ}$ programs, which is quite common in practice. In particular, the main result of this paper is that the convergence rate of linear $\text{\sf Datalog}^{\circ}$ programs under any $p$-stable semiring is $O(pn^{3})$. Next, we study the convergence rate in terms of the number of elements in the semiring for linear $\text{\sf Datalog}^{\circ}$ programs. When $L$ is the number of elements, we show that the convergence rate is bounded by $O(pn\log L)$. This significantly improves the convergence rate for small $L$.
keywords:
Datalog, convergence rate, semiring
1 Introduction
In order to express common recursive computations with aggregates in modern data analytics while retaining the syntactic and semantic simplicity of Datalog, [12] introduced $\text{\sf Datalog}^{\circ}$, an extension of Datalog that allows for aggregation and recursion over an arbitrary commutative partially ordered pre-semiring (POPS). Datalog is exactly $\text{\sf Datalog}^{\circ}$ over the Boolean semiring. Like Datalog, $\text{\sf Datalog}^{\circ}$ has a declarative least fixpoint semantics, and the least fixpoint can be found via the naïve iteration algorithm that repeatedly applies the immediate consequence operator until no further change is possible. Moreover, its additional algebraic structure allows for common recursions with aggregations.
However unlike Datalog, the naïve evaluation of a $\text{\sf Datalog}^{\circ}$ program may not always converge in a finite number of steps. The convergence of a $\text{\sf Datalog}^{\circ}$ program over a given POPS can be studied through its “core semiring”, which is where we focus our attention on in this paper. This paper will only study $\text{\sf Datalog}^{\circ}$ programs over commutative semirings, referring the readers to [12] for the generality of POPS.
It is known that the commutative semirings for which the iterative evaluation of $\text{\sf Datalog}^{\circ}$ programs is guaranteed to converge are exactly those semirings that are stable [12]. A semiring is $p$-stable if the number of iterations required for any one-variable recursive linear $\text{\sf Datalog}^{\circ}$ program to reach a fixed point is at most $p$, and a semiring is stable if there exists a $p$ for which it is $p$-stable. Further, every non-stable semiring has a simple (linear) $\text{\sf Datalog}^{\circ}$ program with one variable with the property that the iterative evaluation of this program over that semiring will not converge. Thus it is natural to concentrate on $\text{\sf Datalog}^{\circ}$ programs over stable semirings. Previously, the best known upper bound on the convergence rate, which is the number of iterations until convergence, is $\sum_{i=1}^{n}(p+2)^{i}=\Theta(p^{n})$ steps, where $n$ is the number of ground atoms for the IDB’s that ever have a nonzero value at some point in the iterative evaluation of the $\text{\sf Datalog}^{\circ}$ program, and the underlying semiring is $p$-stable. In contrast there are no known lower bounds that show that iterative evaluation requires an exponential (in the parameter $n$) number of steps to reach convergence.
The exact general upper bound on the convergence rate of $\text{\sf Datalog}^{\circ}$ programs over $p$-stable semirings is open, even for the special case of linear $\text{\sf Datalog}^{\circ}$ programs. Linear $\text{\sf Datalog}^{\circ}$ programs are quite common in practice. [^1] For example, we can express transitive closure, all-pairs-shortest-paths, or weakly connected components in linear $\text{\sf Datalog}^{\circ}$. Thus, in this paper we focus on this “easiest” case where the exact convergence rate is not known.
Currently, the best known upper bound on the convergence rate of linear $\text{\sf Datalog}^{\circ}$ programs over $p$-stable semirings is $\sum_{i=1}^{n}(p+1)^{i}$ steps. This bound is unsatisfactory in the following sense. The prototypical example of a $p$-stable semiring is the tropical semiring $\text{\sf Trop}_{p}^{+}$, (see Section 2 for a definition of $\text{\sf Trop}_{p}^{+}$); in this case, it is known that the naïve algorithm converges in $O(pn)$ steps for linear $\text{\sf Datalog}^{\circ}$ programs [12]. These results leave open the possibility that the convergence rate of the naïve algorithm on linear $\text{\sf Datalog}^{\circ}$ programs over $p$-stable semirings could be exponentially smaller than the best known guarantee.
This paper seeks to obtain tighter bounds on the convergence rate of naïve evaluation of linear $\text{\sf Datalog}^{\circ}$ programs. As the iterative evaluation of $\text{\sf Datalog}^{\circ}$ programs is a reasonably natural and important algorithm/process, bounding the running time of this process is of both theoretical interests and practical interests. In practice, a known upper bound on the convergence rate allows the database system to determine before hand an upper bound on the number of iterations that will be required to evaluation a particular $\text{\sf Datalog}^{\circ}$ program.
1.1 Background
Before stating our main results, we need to back up to set the stage a bit. A (traditional) Datalog program $\Pi$ consists of a set of rules of the form:
[…]
where $R_{0},\ldots,R_{m}$ are predicate names (not necessarily distinct) and each $\bm{X}_{i}$ is a tuple of variables and/or constants. The atom $R_{0}(\bm{X}_{0})$ is called the head, and the conjunction $R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{m}(\bm{X}_{m})$ is called the body of the rule. Multiple rules with the same head are interpreted as a disjunction. A predicate that occurs in the head of some rule in $\Pi$ is called an intensional database predicate or IDB, otherwise it is called an extensional database predicate or EDB. The EDBs form the input database, and the IDBs represent the output instance computed by $\Pi$. The finite set of all constants occurring in an EDB is called the active domain, and denoted ADom. A Datalog program is linear if every rule has at most one IDB predicate in the body.
There is an implicit existential quantifier over the body for all variables that appear in the body but not in the head, where the domain of the existential quantifier is ADom. In a linear Datalog program every conjunction has at most one IDB.
Example 1.
The textbook example of a linear Datalog program is the following one, which computes the transitive closure of a directed graph, defined by the edge relation $E(X,Y)$:
[…]
Here $E$ is an EDB predicate, $T$ is an IDB predicate, and ADom consists of the vertices in the graph. The other way to write this program is to write it as a union of conjunctive queries (UCQs), where the quantifications are explicit:
[…]
A Datalog program can be thought of as a function (called the immediate consequence operator, or ICO) that maps a set of ground IDB atoms to a set of ground IDB atoms. Every rule in the program is an inference rule that can be used to infer new ground IDB atoms from old ones. For a particular EDB instance, this function has a unique least fixpoint which can be obtained via repeatedly applying the ICO until a fixpoint is reached [1]. This least fixpoint is the semantics of the given Datalog program. The algorithm is called the naïve evaluation algorithm, which converges in a polynomial number of steps in the size of the input database, given that the program is of fixed size.
Like Datalog programs, a $\text{\sf Datalog}^{\circ}$ program consists of a set of rules, where the unions of conjunctive queries are now replaced with sum-product queries over a commutative semiring $\bm{S}=(S,\oplus,\otimes,0,1)$; see Section 2. Each rule has the form:
[…]
where sum ranges over the active ADom of the variables not in $X_{0}$. Further each ground atom in an EDB predicate or IDB predicate is associated with an element of the semiring $\bm{S}$, and the element associated with a tuple in an EDB predicate is specified in the input. The EDBs form the input database, and the ground atoms in the IDB’s that have an associated semiring value that is nonzero represent the output instance computed by the $\text{\sf Datalog}^{\circ}$ program. Note that in a Datalog program the ground atom present in the input or output databases can be thought of as those that are associated with the element $1$ in the standard Boolean semiring. A Datalog program is a $\text{\sf Datalog}^{\circ}$ program over the Boolean semiring $\bm{S}=(\{\text{\sf true},\text{\sf false}\},\vee,\wedge,\text{\sf false},\text{\sf true})$. Again a $\text{\sf Datalog}^{\circ}$ program is linear if every rule has no more than one IDB prediciate in its body.
Example 2.
A simple example of a linear $\text{\sf Datalog}^{\circ}$ program is the following,
[…]
which is (2) with $(\vee,\wedge,\exists_{Z})$ replaced by $(\oplus,\otimes,\bigoplus_{Z})$.
When interpreted over the Boolean semiring, we obtain the transitive closure program from Example 1. When interpreted over the tropical semiring $\text{\sf Trop}^{+}=(\mathbb{R}_{+}\cup\{\infty\},\min,+,\infty,0)$, we have the All-Pairs-Shortest-Path (APSP) program, which computes the shortest path length $T(X,Y)$ between all pairs $X,Y$ of vertices in a directed graph specified by an edge relation $E(X,Y)$, where the semiring element associated with $E(X,Y)$ is the length of the directed edge $(X,Y)$ in the graph:
[…]
A $\text{\sf Datalog}^{\circ}$ program can be thought of as an immediate consequence operator (ICO). To understand how the ICO works in $\text{\sf Datalog}^{\circ}$, consider a rule with head $R$ and let $A$ be a ground atom for $R$ with associated semiring value $y$, and assume that for $A$ the body of this rule evaluates to the semiring element $x$. As a result of this, the ICO associates $A$ with with $x\oplus y$. Note that the functioning of the $\text{\sf Datalog}^{\circ}$ ICO, when the semiring is the standard Boolean semiring, is identical to how the Datalog ICO functions. The iterative evaluation of a $\text{\sf Datalog}^{\circ}$ program works in rounds/steps, where initially the semiring element 0 is associated with each possible ground atom in an IDB, and on each round the ICO is applied to the current state. So in the context of the $\text{\sf Datalog}^{\circ}$ program in Example 1, if $(a,b)$ was a ground atom in $T$ with associated semiring element $x$, meaning that the shortest known directed path from vertex $a$ to vertex $b$ has length $x$, and $(b,c)$ was a ground atom in $E$ with associated semiring element $y$, meaning that there is a directed edge from $b$ to $c$ with length $y$, then the ICO would make the semiring element associated with $A$ the minimum (as minimum is the addition operation in the tropical semiring) of its current value and $x+y$ (as normal additional is the multiplication operation in the tropical semiring).
Since the final associated semiring values of the ground atoms in an IDB are not initially known, it is natural to think of them as (IDB) variables. Then the grounded version of the ICO of a $\text{\sf Datalog}^{\circ}$program is a map $\bm{f}:S^{n}\rightarrow S^{n}$, where $n$ is the number of ground atoms for the IDB’s that ever have a nonzero value at some point in the iterative evaluation of the $\text{\sf Datalog}^{\circ}$ program. For instance, in (5), there would be one variable for each pair $(x,y)$ of vertices where there is a directed path from $x$ to $y$ in the graph. So the grounded version of the ICO of a $\text{\sf Datalog}^{\circ}$ program has the following form:
$-\displaystyle X_{1}\text{ :- }$- $f_{1}(X_{1},\ldots,X_{n})$
$\ldots$ (6)
$-\displaystyle X_{n}\text{ :- }$- $f_{n}(X_{1},\ldots,X_{n})$
where the $X_{i}$’s are the IDB variables, and $f_{i}$ is the component of $\bm{f}$ corresponding to the IDB variable $X_{i}$. Note that each component function $f_{i}$ is a multivariate polynomial in the IBD variables of degree at most the maximum number of terms in any product in the body of some rule in the $\text{\sf Datalog}^{\circ}$ program. After $q$ iterations of the iterative evaluation of a $\text{\sf Datalog}^{\circ}$ program, the semiring value associated with the ground atom corresponding to $X_{i}$ will be:
${f}_{i}^{(q)}(\bm{0})$ (7)
Definition 1.1 ($p$-stability).
Given a semiring $\bm{S}=(S,\oplus,\otimes,0,1)$ and $u\in S$, let $u^{(p)}:=1\oplus u\oplus u^{2}\oplus\cdots\oplus u^{p}$, where $u^{i}:=u\otimes u\otimes\cdots\otimes u$ ($i$ times). Then $u\in S$ is $p$-stable if $u^{(p)}=u^{(p+1)}$, and semiring $\bm{S}$ is $p$-stable if every element $u\in S$ is $p$-stable.
Definition 1.2 (Stability index and convergence rate).
A function $f:S^{n}\to S^{n}$ is $p$-stable if $f^{(p+1)}(\bm{0})=f^{(p)}(\bm{0})$, where $f^{(k)}$ is the $k$-fold composition of $f$ with itself. The stability index of $f$ is the smallest $p$ such that $f$ is $p$-stable. The convergence rate of a $\text{\sf Datalog}^{\circ}$ program is the stability index of its ICO.
The following bounds on the rate of convergence of a general (multivariate) polynomial function $f:S^{n}\to S^{n}$, where $\bm{S}$ is $p$-stable; this result naturally infers bounds on the convergence rate of $\text{\sf Datalog}^{\circ}$ programs over $p$-stable semirings.
Theorem 1.3 ([12]).
The convergence rate of a $\text{\sf Datalog}^{\circ}$ program over a $p$-stable commutative semiring is at most $\sum_{i=1}^{n}(p+2)^{i}$. Further, the convergence rate is at most $\sum_{i=1}^{n}(p+1)^{i}$ if the $\text{\sf Datalog}^{\circ}$ program is linear. Finally, the convergence rate of a $\text{\sf Datalog}^{\circ}$ program is at most $n$ if $p=0$.
Thus the natural question left open by  [12] was whether these upper bounds on the rate of convergence are (approximately) tight, and thus convergence can be exponential, or whether significantly better bounds are achievable.
1.2 Our Results
When a $\text{\sf Datalog}^{\circ}$ program over a semiring $\bm{S}$ is linear, its ICO $f:S^{n}\rightarrow S^{n}$ is a linear function of the form $f(x)=A\otimes x\oplus b$, where $A$ is an $n$ by $n$ matrix with entries from $S$, $b$ is an $n\times 1$ column vector with entries from $S$, and the scalar multiplication and addition are from $\bm{S}$. To simplify notations, we will use $+$ and $\cdot$ to denote the operations $\oplus$ and $\otimes$ respectively. Further, following the convention, we may omit $\otimes$ if it is clear from the context. Then, we have
$f(x)=Ax+b$
Example 3.
For the APSP $\text{\sf Datalog}^{\circ}$ program, $n$ is the number of edges in the graph, $b_{(u,v)}$ is $E(u,v)$, and the matrix $A$ would be:
$A_{(u,v),(u,w)}=\begin{cases}E(v,w)&\text{if }u\neq v\\ 0&\text{otherwise}\end{cases}$ (8)
The stability index of $f$ can easily be expressed in terms of the matrix $A$ and vector $b$. Letting $A^{0}=I$ where $I$ is the identity matrix, we have
$f^{(k)}(x)=A^{k}x+A^{(k-1)}b$ $\text{where }A^{(k)}:=\sum_{h=0}^{k-1}A^{h}.$
Thus, the linear function $f=Ax+b$ is $p$-stable if and only if $A^{(p)}b=A^{(p+1)}b$. Using the simple form of $f$ for the linear case, we can rewrite Definition 1.2 into the following simpler form. {observation}[Convergence Rate of a Linear $\text{\sf Datalog}^{\circ}$ Program]
  • The convergence rate for a particular linear $\text{\sf Datalog}^{\circ}$ program, with associated matrix $A$ and vector $b$, is the minimum natural number $p$ such that $A^{(p)}b=A^{(p+1)}b$.
  • The convergence rate for general linear $\text{\sf Datalog}^{\circ}$ programs over a semiring $S$ is then the maximum over all choices of $A$ and $b$, of the convergence rate for that particular $A$ and $b$.
Our first result is a bound of $O(pn^{3})$ on the rate of convergence for linear $\text{\sf Datalog}^{\circ}$ programs.
Theorem 1.4.
Every linear $\text{\sf Datalog}^{\circ}$ program over a $p$-stable commutative semiring $\bm{S}$ converges in $O(pn^{3})$ steps, where $n$ is the total number of ground IDB atoms.
The proof Theorem 1.4 can be found in Section 3, but we will give a brief overview of the main ideas of the proof here. Consider the complete $n$-vertex loop-digraph $G$ where the edge from $i$ to $j$ is labeled with entry $A_{i,j}$. Then note that row $i$ column $j$ of $A^{(h)}$ is the sum – over all walks $W$ from $i$ to $j$ of length $\leq h$ – of the product of the edge labels on the walk. We show that the summand corresponding to a walk $W$ with more than $h=\Omega(pn^{3})$ hops doesn’t change the sum $A^{(h)}_{i,j}$. We accomplish this by rewriting the summand corresponding to $W$, using the commutativity of multiplication in $S$, as the product of a simple path and of multiple copies of at most $n^{2}-n$ distinct closed walks. We then note that, by the pigeon hole principle, one of these closed walks, say $C$, must have mulitplicity greater than $p$. We then conclude that the summand corresponding to $W$ will not change the sum $A^{(h)}_{i,j}$ by appealing to the stability of the semiring element that is the product of the edges in $C$.
In the conference version of this paper [10], and the original arxiv version of the paper, we claimed a matching lower bound. That is we claimed that for any $p,n\geq 1$, there is a linear $\text{\sf Datalog}^{\circ}$ program over a $p$-stable semiring $\bm{S}$ that requires $\Omega(pn^{3})$ steps to converge. It was subsequently pointed out that our claim was incorrect [6]. Still in Section 4 we briefly discuss our lower bound attempt, and the underlying conceptual error that led to our mistake, as we believe that it is instructive.
A natural question is whether $\text{\sf Datalog}^{\circ}$ programs over semirings with small sets will converge more quickly. In section 5 we answer this question in the affirmative by showing that the rate of convergence of a $\text{\sf Datalog}^{\circ}$ program over a $p$-stable commutative semiring with $L$ elements is $O(pn\log L)$.
Theorem 1.5.
Every linear $\text{\sf Datalog}^{\circ}$ program over a $p$-stable commutative semiring that contains $L$ elements in its ground set converges in $O(pn\log L)$ steps.
Let us explain the high level idea of the proof, with some simplifying assumptions, starting with the assumption that the stability of $\bm{S}$ is $p=1$. Again we think of $A^{(k)}_{i,j}$ as the sum of products over walks $W$ from $i$ to $j$ with at most $k$ hops. Now consider a walk $W$ from $i$ to $j$ consisting of $\Omega(n\log L)$ hops. Since there are $n$ vertices, there exists a vertex $v$ such that there are at least $\Omega(\log L)$ prefixes $P_{1},P_{2},\ldots,P_{q}$ of $W$ ending at $v$. Let $C_{i}$ be the closed walk starting and ending at $v$ such that appending $C_{i}$ to $P_{i}$ produces $P_{i+1}$. Let us make the simplifying assumption that all of these closed walks $C_{i}$ are distinct. The key observation is that if we delete any subset $\mathcal{C}$ of these $\Omega(\log L)$ closed walks from $W$, the result will still be a walk from $i$ to $j$. Thus there are at least $2^{\Omega(\log L)}$ walks from $i$ to $j$ that can be formed by deleting a subset $\mathcal{C}$ of these closed walk. Since there are at most $L$ distinct elements, by the pigeon hole principle, there must be some element $e$ of $S$ such that that are two subsets $\mathcal{C},\mathcal{C}^{\prime}$ where the product of the edges in them are the same. We then conclude by the stability of $e$ that the summand corresponding to $W$ does not change the sum $A^{(k)}_{i,j}$.
In the conference version of this paper [10], and the original arxiv version of the paper, we claimed a nearly matching lower bound. We now believe that this lower bound is also not correct, for reasons roughly similar to those that led to our incorrect $\Omega(pn^{3})$ lower bound.
Finally in the appendix we consider a special case from [12], in which the commutative semiring $\bm{S}$ is naturally ordered.
2 Preliminaries
In this section, we introduce the notation and terminology used throughout the paper.
Definition 2.1 (Semiring).
A semiring [8] is a tuple $\bm{S}=(S,\oplus,\otimes,0,1)$ where
  • $\oplus$ and $\otimes$ are binary operators on $S$,
  • $(S,\oplus,0)$ is a commutative monoid, meaning $\oplus$ is commutative and associative, and 00 is the identity for $\oplus$,
  • $(S,\otimes,1)$ is a monoid, meaning $\otimes$ is associative, and 00 is the identity for $\oplus$,
  • 00 annilates every element $a\in S$, that is $a\otimes 0=0\otimes a=0$, and
  • $\otimes$ distributes over $\oplus$.
A commutative semiring $\bm{S}=(S,\oplus,\otimes,0,1)$ is a semiring where additionally $\otimes$ is commutative.
If $A$ is a set and $p\geq 0$ a natural number, then we denote by $\mathcal{B}_{p}(A)$ the set of bags (multiset) of $A$ of size $p$, and $\mathcal{B}_{\texttt{fin}}(A):=\bigcup_{p\geq 0}\mathcal{B}_{p}(A)$. We denote bags as in $\{\{a,a,a,b,c,c\}\}$. Given $\bm{x},\bm{y}\in\mathcal{B}_{fin}(\mathbb{R}_{+}\cup\infty)$, define
$\bm{x}\uplus\bm{y}$ $:=\mbox{bag union of $\bm{x},\bm{y}$}$ $\bm{x}+\bm{y}$ :=$:=\{\{{u+v}\mid{u\in\bm{x},v\in\bm{y}}\}\}$
Example 4.
For any multiset $\bm{x}=\{\{x_{0},x_{1},\ldots,x_{n}\}\}$, where $x_{0}\leq x_{1}\leq\ldots\leq x_{n}$, and any $p\geq 0$, define:
${\min}_{p}(\bm{x})$ :=$:=\{\{x_{0},x_{1},\ldots,x_{\min(p,n)}\}\}$
In other words, $\min_{p}$ returns the smallest $p+1$ elements of the bag $\bm{x}$. Then, for any $p\geq 0$, the following is a semiring:
$\text{\sf Trop}^{+}_{p}$ $:=(\mathcal{B}_{p+1}(\mathbb{R}_{+}\cup\{\infty\}),\oplus_{p},\otimes_{p},\bm{0}_{p},\bm{1}_{p})$
where:
[…]
$\bm{x}\otimes_{p}\bm{y}\stackrel{{\scriptstyle\text{def}}}{{=}}$ ${\min}_{p}(\bm{x}+\bm{y})$ $\bm{1}_{p}\stackrel{{\scriptstyle\text{def}}}{{=}}$$\{0,\infty,\ldots,\infty\}$
For example, if $p=2$ then $\{\{3,7,9\}\}\oplus_{2}\{\{3,7,7\}\}=\{\{3,3,7\}\}$ and $\{\{3,7,9\}\}\otimes_{2}\{\{3,7,7\}\}=\{\{6,10,10\}\}$. The following identities are easily checked, for any two finite bags $\bm{x},\bm{y}$:
${\min}_{p}({\min}_{p}(\bm{x})\uplus{\min}_{p}(\bm{y}))=$ ${\min}_{p}(\bm{x}\uplus\bm{y})$ ${\min}_{p}({\min}_{p}(\bm{x})+{\min}_{p}(\bm{y}))=$ ${\min}_{p}(\bm{x}+\bm{y})$ (9)
Note that $\text{\sf Trop}^{+}_{0}$ is the natural “min-plus” semiring that we used in Example 2.
In the following fact about $p$-stable commutative semiring will be useful.
Proposition 2.2.
Given a $p$-stable commutative semiring $\bm{S}=(S,\oplus,\otimes,0,1)$, for any $u\in S$, we have $pu=(p+1)u$, where $pu$ here is shorthand for $\oplus_{i=1}^{p}u$.
Proof 2.3.
This follows directly from the $p$-stability of 1, and the fact that $11^{2}=$ 1.
Let us now explain the general procedure for creating a matrix $A$ and a vector $b$ from a linear $\text{\sf Datalog}^{\circ}$ program $Q$. Each ground tuple for each IDB predicate $R$ can be viewed as a variable, and ground tuples of EDB predicates can be viewed as constants. Then a $\text{\sf Datalog}^{\circ}$ rule, where the head is IDB predicate $R$, can be converted into a collection of linear equations, one for each ground tuble of $R$. Since all rules that share IDB $R$ as a the head can be combined via $\oplus$, we can compactly rewrite the entire set of $\text{\sf Datalog}^{\circ}$ rules as a collection of linear equations of the following form:
$T_{i}\leftarrow\bigoplus_{j=1}^{n}(A_{ij}\otimes T_{j})\oplus b_{i}$
where $T_{1},T_{2},\ldots,T_{n}$ are the variables corresponding to ground tuples of IDB predicates. By using more familiar notation $+$, $\cdot$ in the lieu of $\otimes$,$\oplus$ , ⊕, we can model a linear $\text{\sf Datalog}^{\circ}$ program $Q$ by a linear function $f:S^{n}\rightarrow S^{n}$ of the form:
$f(x)=Ax+b$
where $n$ is the total number of ground tuples across all IDB predicates, $x$ is an $n$-dimensional vector with entries from $S$, $A$ is an $n$ by $n$ matrix with entries from the $S$, and $b$ is a dimensional column vector with entries from $S$. For some more examples of converting linear $\text{\sf Datalog}^{\circ}$ programs into linear equations see  [12].
3 Upper bounding the Convergence as a Function of the Matrix Dimension and the Semiring Stability
This section is devoted to proving Theorem 1.4, the main theorem of the paper. More specifically, we will show the following lemma. This lemma upper bounds the convergence of a $p$-stable semiring and implies Theorem 1.4.
Lemma 3.1.
Let $A$ be an $n\times n$ matrix over a $p$-stable semiring $S$. Then $A^{(k+1)}=A^{(k)}$, where $k=n(n^{2}-n)(p+2)+n-1$.
Consider the complete $n$-vertex loop-digraph $G$ where the edge from $i$ to $j$ is labeled with entry $A_{i,j}$. Then
$A^{h}_{i,j}=\sum_{W\in{\mathcal{W}^{h}_{i,j}}}\Phi(W)$
where ${\mathcal{W}^{h}_{i,j}}$ is the collection of all $h$-hop walks from $i$ to $j$ in $G$, and
$\Phi(W)=\prod_{(a,b)\in W}A_{a,b}$
is the product off all the labels on all the directed edges in $W$. That is, row $i$ column $j$ of $A^{h}$ is the sum over all $h$-hop walks $W$ from $i$ to $j$ of the product of the labels on the walk. Similarly then,
$A^{(h)}_{i,j}=\sum_{g=0}^{h}A^{g}_{i,j}=\sum_{g=0}^{h}\sum_{W\in{\mathcal{W}^{g}_{i,j}}}\Phi(W)$
That is, row $i$ column $j$ of $A^{(h)}$ the sum over all walks $W$ from $i$ to $j$ with at most $h$ hops of the product of the labels on the walk. Further,
$A^{(h+1)}_{i,j}=A^{(h)}_{i,j}+A^{h+1}_{i,j}=A^{(h)}_{i,j}+\sum_{W\in{\mathcal{W}^{h+1}_{i,j}}}\Phi(W)$
Our proof technique is to show that, by the $p$-stability of $S$, it must be the case that for each $W\in{\mathcal{W}^{p+1}_{i,j}}$ it is the case that
$A^{(k)}_{i,j}+\Phi(W)=A^{(k)}_{i,j}$
The proof that $A^{(k)}_{i,j}=A^{(k+1)}_{i,j}$ then immediately follows by applying this fact to each $W\in{\mathcal{W}^{k+1}_{i,j}}$.
Fix $W=i,\dots,j$ to be an arbitrary walk in ${\mathcal{W}^{k+1}_{i,j}}$. Our next intermediate goal is to rewrite $\Phi(W)$ using the commutativity of multiplication in $S$ as the product of a simple path and at most $n^{2}-n$ multicycles. That is,
$\Phi(W)=\Phi(P)\prod_{h=1}^{\ell}\Phi(C_{h}^{z_{h}})$
where $P$ is a simple path from $i$ to $j$ in $G$, each $C_{h}$ is a simple cycle in $W$ that is repeated $z_{h}$ times, and $\ell\leq n^{2}-n$. We accomplish this goal via the following tail-recursive construction. The recursion is passed a collection of edges, and a parameter $h$. Initially, the edges are those in $W$ and $h$ is set to zero.
Recursive Construction: The base case is if $W$ is a simple path. In the base case the path $P$ is set to $W$ and $\ell$ is set to 00. Otherwise:
  • $h$ is incremented
  • $C_{h}$ is set to be an arbitrary simple cycle in $W$.
  • Let $z_{h}$ be the minimum over all edges $e\in C_{h}$ of the number of times that $e$ is traversed in $W$.
  • Let $W^{\prime}$ be the collection of edges in $W$ except that $z_{h}$ copies of every edge in $C_{h}$ are removed.
  • The construction then recurses on $W^{\prime}$ and $h$.
In Lemma 3.3 we show that a particular statement about $W$ is invariant through the recursive construction. A proof Lemma 3.3 requires the following lemma. The proof of the following lemma (or at least, the techniques needed for a proof) can be found in most introductory graph theory texts, e.g. [22] Theorem 23.1.
Lemma 3.2.
  • A loop digraph $G$ has a Eulerian walk from from a vertex $i$ to a vertex $j$, where $i\neq j$, if and only if vertex $i$ has out-degree one greater than its in-degree, vertex $j$ has out-degree one less than its in-degree, every other vertex has equal in-degree and out-degree, and all of the vertices with nonzero degree belong to a single connected component of the underlying undirected graph.
  • A loop digraph has an Eulerian cycle that includes a vertex $i$ if and only if every vertex has equal in-degree and out-degree, vertex $i$ has non-zero in-degree, and all of the vertices with nonzero degree belong to a single connected component of the underlying undirected graph.
Lemma 3.3.
Let $W$ be a collection of edges that is passed at some point in the recursive construction. Let $D_{1},\ldots,D_{h}$ be a partition of the edges of $W$ with the property that if the edges in $W$ were viewed as undirected, then the connected components would be $D_{1},\ldots,D_{h}$.
  • If $i\in D_{f}$ then $D_{f}$ is a walk from $i$ to $j$.
  • If $i\notin D_{f}$ then $D_{f}$ is a Eulearian circuit.
Proof 3.4.
The proof is by induction on the number of steps of the recursive construction. The statement is obviously true for the initial walk $W$, which is the base case. Now consider one step of the recursive construction. Removing copies of a cycle from $W$ does not change the difference between the in-degree and out-degree of any vertex. Thus by Lemma 3.2 the only issue we need to consider is vertex $i$ and vertex $j$ possibly ending up in different connected components of $W^{\prime}$. Let $D_{f}$ be the connected component of $W$ that contains $i$ (which also contains $j$ by the induction hypothesis), let $D^{\prime}_{a}$ be the connected component of $W^{\prime}$ that contains $i$, and let $D^{\prime}_{b}$ be the connected component of $W^{\prime}$ that contains $j$, where $a\neq b$. Then the walk in $D_{f}$ from $i$ to $j$ must cross the cut (in either direction) formed by the vertices in $D^{\prime}_{a}$ an odd number of times. But the edges in $C_{h}$ must cross the cut (in either direction) formed by the vertices in $D^{\prime}_{a}$ an even number of times. Thus there must be an edge in $D_{f}$ minus $z_{h}$ copies of $C_{h}$ that must cross the cut (in either direction) formed by the vertices in $D^{\prime}_{a}$. However, then this is a contradiction to $D^{\prime}_{a}$ be a connected component in $W^{\prime}$.
We now make a sequence of observations, that will eventually lead us to our proof of Lemma 3.1.
{observation}
  • $1\leq\ell\leq n^{2}-n$.
  • The recursive construction terminates.
  • $\Phi(W)=\Phi(P)\prod_{h=1}^{\ell}\Phi(C_{h}^{z_{h}})$.
Proof 3.5.
As $W$ contains $k+1=n(n^{2}-n)(p+2)+n$ edges, then it must contain a simple cycle, which implies $\ell\geq 1$. Consider one iteration of our recursive construction. There must be a directed edge $e\in C_{h}$ that appears exactly $z_{h}$ times in $W$. Thus there are no occurrences of $e$ in $W^{\prime}$. Thus $e$ can not appear in any future cycles, that is $e\notin C_{g}$ for any $g>h$. The first observation then follows because there are at most $n^{2}-n$ different edges in $G$. The second observation is then an immediate consequence of the first observation, and the invariant established in Lemma 3.3 . The third observation follows because no edges are ever lost or created in the recursive construction.
{observation}
There is a cycle $C_{s}$, $1\leq s\leq\ell$, such that $z_{s}$ is at least $p+2$.
Proof 3.6.
Since $P$ is a simple path it contains at most $n-1$ edges. Thus $W-P$ contains at least $n(n^{2}-n)(p+2)$ edges. As any simple cycle contains at most $n$ edges, and as there are at most $\ell\leq n^{2}-n$ cycles, then by applying the pigeon hole principle to the cycle decomposition of $W$ we can conclude that there must be a cycle $C_{s}$ that has multiplicity at least $p+2$, that is $z_{s}\geq p+2$.
For convenience, consider renumbering the cycles so that $s=1$ where $z_{s}\geq p+1$, which exists by Observation 3. For each $h$ such that $11\leq h\leq p+$ 1, let define $W_{h}$ be the collection of edges in $W$ minus $h$ copies of every edge in $C_{1}$.
{observation}
The edges in each $W_{h}$, $11\leq h\leq p+$ 1 form a walk from $i$ to $j$.
Proof 3.7.
As $z_{1}\geq p+2$ and $h\leq p+1$, every edge that appears in $W$ also appears in $W_{h}$. So $W_{h}$ has the same connectivity properties as $W$, and $W_{h}$ has the same vertices with positive in-degree as does $W$. Also as $C_{1}$ is a simple cycle, the difference between in-degree and out-degree for each vertex is the same in $W_{h}$ as in $W$. Thus the result follows by appealing to Lemma 3.2.
{observation}
For each $h$ such that $11\leq h\leq p+$ 1 it is the case that
$\Phi(W_{h})=\Phi(P)\Phi(C_{1}^{z_{1}-h})\prod_{f=2}^{\ell}\Phi(C_{f}^{z_{f}})$
Proof 3.8.
This follows directly from the defintion of $W_{h}$.
{observation}
For all $h$ such that $11\leq h\leq p+$ 1, we have $W_{h}\in\bigcup_{f=0}^{k}{\mathcal{W}^{f}_{i,j}}$. That is $\Phi(W_{h})$ appears as a term in $A_{i,j}^{(k)}$.
Proof 3.9.
This follows because $W$ has $k+1$ edges and $C_{1}$ is non-empty, so removing edges in $C_{i}$ strictly decreases the number of edges.
We are now ready to prove Lemma 3.1. By Observation 3 we know that each $\Phi(W_{h})$ is included in $A^{(k)}$, and thus there exists an element $r$ in the semiring $S$ such that
$A^{(k)}_{i,j}=r+\sum_{h=1}^{p+1}\Phi(W_{h})$
Thus
$A^{(k)}_{i,j}+\Phi(W)$ $=r+\sum_{h=1}^{p+1}\Phi(W_{h})+\Phi(W)$
$=r+\sum_{h=1}^{p+1}\left(\Phi(P)\Phi(C_{1}^{z_{1}-h})\prod_{f=2}^{\ell}\Phi(C_{f}^{z_{f}})\right)+\Phi(P)\prod_{f=1}^{\ell}\Phi(C_{f}^{z_{f}})$
$=r+\left(\Phi(P)\prod_{f=2}^{\ell}\Phi(C_{f}^{z_{f}})\right)\left(\sum_{h=1}^{p+1}\Phi(C_{1}^{z_{1}-h})+\Phi(C_{1}^{z_{1}})\right)$ (10)
$=r+\left(\Phi(P)\prod_{f=2}^{\ell}\Phi(C_{f}^{z_{f}})\right)\left(\Phi(C_{1}^{z_{1}-(p+1)})\sum_{h=0}^{p+1}\Phi(C_{1}^{h})\right)$
$=r+\left(\Phi(P)\prod_{f=2}^{\ell}\Phi(C_{f}^{z_{f}})\right)\left(\Phi(C_{1}^{z_{1}-(p+1)})\sum_{h=0}^{p+1}\left[\Phi(C_{1})\right]^{h}\right)$ (11)
$=r+\left(\Phi(P)\prod_{f=2}^{\ell}\Phi(C_{f}^{z_{f}})\right)\left(\Phi(C_{1}^{z_{1}-(p+1)})\sum_{h=0}^{p}\left[\Phi(C_{1})\right]^{h}\right)$ (12)
$=r+\left(\Phi(P)\prod_{f=2}^{\ell}\Phi(C_{f}^{z_{f}})\right)\left(\Phi(C_{1}^{z_{1}-(p+1)})\sum_{h=0}^{p}\Phi(C_{1}^{h})\right)$
$=r+\left(\Phi(P)\prod_{f=2}^{\ell}\Phi(C_{f}^{z_{f}})\right)\left(\sum_{h=1}^{p+1}\Phi(C_{1}^{z_{1}-h})\right)$
$=r+\sum_{h=1}^{p+1}\left(\Phi(P)\Phi(C_{1}^{z_{1}-h})\prod_{f=2}^{\ell}\Phi(C_{f}^{z_{f}})\right)$
$=r+\sum_{h=1}^{p+1}\Phi(W_{h})=A^{(k)}_{i,j}$
The equality in line (10) follows from Observation 3. The equality in line (11) follows from the defintion of $\Phi$. The key step in this line of equations is the equality in line (12), which follows from the stability of $\Phi(C_{1})$. The rest of the equalities follow from basic algebraic properties of semirings, or by definition of the relavent term.
4 A Failed Lower Bound Attempt for Convergence Rate of Linear $\text{\sf Datalog}^{\circ}$ Programs
Recall that the definition of $p$-stability is that
$1\oplus u\oplus u^{2}\oplus\cdots\oplus u^{p}=1\oplus u\oplus u^{2}\oplus\cdots\oplus u^{p}\oplus u^{p+1}$
Thus it is natural (but wrong) to conjecture that in order for a walk $W$ to be absorbed by shorter walks that it must be the case that one can express $\Phi(W)$ as $XY^{p+1}$ where $X$ and $Y^{p+1}$ are products of two parts of some partition of $W$. In this case then there is at least the possibility that $W$ is absorbed by
$X\oplus X\otimes Y\oplus X\otimes Y^{2}\ldots\oplus X\otimes Y^{p}$
if it happened that there were shorter walks with products equal to these summands. We now demonstrate that in order to guarantee that a walk $W$ has the property that it can be decomposed so that $\Phi(W)=X\otimes Y^{p+1}$ then $W$ has to be of length $\Omega(pn^{3})$.
We define a graph $G$. Let $G$ be a directed graph where the vertex set is the integers from 1 to $n$ inclusive. For simplicity assume $n$ is divisible by 3. The vertices are partitioned into 3 parts, $B=\{1,\ldots,n/3\}$, $C=\{n/3+1,\ldots,2n/3\}$ and $D=\{2n/3+1,\ldots n\}$. There is a directed edge from each vertex in $B$ to vertex $n/3+1$, there is a directed edge from vertex $32n/$ 3 to each vertex in $D$, and there is a directed edge from each vertex in $D$ to each vertex in $B$. Finally, all vertices in $C$ are sequentially connected from $n/3+1$ to $32n/$ 3, i.e., there is a directed edge from $\tau$ to $\tau+1$ for all $\tau\in[n/3+1,2n/3-1]$. See Figure 1.
Figure 1: The graph $G$.
Note that there are $|B|\cdot|D|=n^{2}/9$ edges from $D$ to $B$. Consider a long walk, say from $i:=n/3+1$ to $j:=n/3+2$, that visits each edge from $D$ to $B$ exactly $p$ times. Clearly, for such a walk $W$, we cannot express $W$ as $XY^{p+1}$ for some $X$ and $Y$. Observe that the walk must visit all edges within $C$ (and the edge from $32n/$ 3 to $12n/3+$ 1) before visiting exactly one edge from $D$ to $B$. Thus, to visit each edge from $D$ to $B$ at least $p$ times, the walk must have length at least $p|B|\cdot|D|\cdot|C|=pn^{3}/27$.
5 Bounding Convergence in Terms of the Semiring Ground Set Size
This section investigates the convergence rate with the assumption that the semiring has a ground set of size at most $L$. With this assumption, we can prove significantly better upper bounds. This section’s goal is to prove Theorem 1.5. We prove the following lemma, which will immediately imply Theorem 1.5. As before, we let $\Phi(W)=\prod_{e\in W}e$ for a walk $W$.
Lemma 5.1.
Let $A$ be an $n$ by $n$ matrix both a $p$-stable semiring $S$ with a ground set consisting of $L$ elements. Then $A^{(k+1)}=A^{(k)}$, where $k=\lceil 8p(\lg L+1)n\rceil+1=O(np\lg L)$.
Proof 5.2.
Fix $i,j\in[n]$. Consider any $W\in\mathcal{W}_{i,j}^{k+1}$ for the value of $k$ stated in the lemma. To show the lemma it suffices to show $\Phi(W)+A_{i,j}^{(k)}=A_{i,j}^{(k)}$. By the pigeon hole principle, there must exist a vertex $v$ visited at least $18p(\lg L+1)+$ 1 times. We now conceptually cut $W$ at visits to $v$ to form a cycle decomposition of $W$. That is, we can write $W$ as $TC_{1},\ldots,C_{h}T^{\prime}$ where $T$ is a walk from $i$ to $v$ such that the only time it visits vertex $v$ is on the last step, each $C_{i}$ is a closed walk that includes vertex $v$ exactly once, and $T^{\prime}$ is a walk from $v$ to $j$ that doesn’t visit $v$ again after initially leaving $v$. Note by the definition of vertex $v$ it must be the case that $h\geq 8p(\lg L+1)$. Let $\mathcal{C}=\{C_{f}\mid 1\leq f\leq h\}$ be the collection of cycles in this cycle decomposition.
We now partition $\mathcal{C}$ into parts ${\mathcal{C}}_{1},\ldots,{\mathcal{C}}_{\ell},{\mathcal{J}}$ where for each closed walk $C$ that has multiplicity $m$ in $\mathcal{C}$ there are $2^{f}$ copies of $C$ in $\mathcal{C}_{f}$ and $m-2^{f}$ copies of $C$ in $\mathcal{J}$ where $f=\lfloor\lg m\rfloor$. For a collection $\mathcal{D}$ of cycles, it will be convenient to use $\Phi({\mathcal{D}})$ to denote $\Phi(\bigcup{\mathcal{D}})$. Thus it then immediate that $\Phi(W)=\Phi(T)\Phi(\mathcal{C})\Phi(T^{\prime})\Phi({\mathcal{J}})$. Note that the cardinality of the multiset $\bigcup_{f=1}^{\ell}\mathcal{C}_{f}$ is at least $4p(\lg L+1)$.
We change $\mathcal{C}$ repeatedly without changing $\Phi(\mathcal{C})$. When changing $\mathcal{C}$ into $\mathcal{C}^{\prime}$, we satisfy:
  1. 1.
    $\Phi(\mathcal{C})=\Phi(\mathcal{C}^{\prime})$.
  2. 2.
    $N(\mathcal{C})=N(\mathcal{C}^{\prime})$, where $N(\mathcal{C})$ denotes the size of multi-set $\mathcal{C}$. So, a cycle $C$ contributes to $N(\mathcal{C})$ by the number of times it appears in $\mathcal{C}$. Further, $\mathcal{C}^{\prime}$ has no more edges in total than $\mathcal{C}$ when we count 1 for each edge in one cycle appearance, i.e., $\sum_{C^{\prime}\in\mathcal{C}^{\prime}}|C^{\prime}|\leq\sum_{C\in\mathcal{C}}|C|$.
  3. 3.
    Consider $\langle\ldots,N_{3}(\mathcal{C}^{\prime}),N_{2}(\mathcal{C}^{\prime}),N_{1}(\mathcal{C}^{\prime})\rangle$ and $\langle\ldots,N_{3}(\mathcal{C}),N_{2}(\mathcal{C}),N_{1}(\mathcal{C})\rangle$. The first vector dominates the second lexicographically. Here $N_{\ell}(\mathcal{C})$ denotes the number of cycles in $\mathcal{C}$ of exponent $2^{\ell}$.
  4. 4.
    When we terminate, $N_{\ell}(\mathcal{C})\leq 2\lfloor\lg L+1\rfloor$ for all $\ell\leq\lfloor\lg p\rfloor$.
  5. 5.
    Every cycle in $\mathcal{C}^{\prime}$ also appears in $\mathcal{C}$.
We now describe the transformation from $\mathcal{C}$ into $\mathcal{C}^{\prime}$. Consider the smallest $\ell$ such that $N_{\ell}>2\lfloor\lg L+1\rfloor$. Consider every subset of cardinality $\lfloor\lg L+1\rfloor$, that consists of cycles of exponent $2^{\ell}$; so they are in $\mathcal{C}_{\ell}$. The number of such subsets is at least (${2\lfloor\lg L+1\rfloor\choose\lfloor\lg L+1\rfloor}>2^{\lg L}=L$. Since the semiring has at most $L$ distinct elements, there must exist distinct subsets $A$ and $B$ of cycles of exponent $2^{\ell}$ such that $|A|=|B|$ and $\Phi(A)=\Phi(B)$. Assume wlog that $B$’s cycles have no more edges in total than $A$’s cycles. Then, we replace $A$ with $B$ in $\mathcal{C}$ and let $\mathcal{C}^{\prime}$ be $\mathcal{C}$ after this change.
It is easy to see that $\Phi(\mathcal{C})=\Phi(\mathcal{C}^{\prime})$. This is because we replaced $A$ with $B$ such that $\Phi(A)=\Phi(B)$. More formally, $\Phi(\mathcal{C})=\prod_{\ell^{\prime}}\Phi(\mathcal{C}_{\ell^{\prime}})=(\Phi(\mathcal{C}_{\ell}\setminus A)\Phi(A))\prod_{\ell^{\prime}\neq\ell}\Phi(\mathcal{C}_{\ell^{\prime}})=(\Phi(\mathcal{C}_{\ell}\setminus A)\Phi(B))\prod_{\ell^{\prime}\neq\ell}\Phi(\mathcal{C}_{\ell^{\prime}})=\Phi(\mathcal{C}^{\prime})$. The second property is also obvious because when we replace $A$ with $B$ in the change, we ensured $|A|=|B|$, which implies that the multiset size remains unchanged. Also, it is immediate that the total number of edges don’t increase as $B$ has no more edges in total than $A$. The forth property, the termination condition, is immediate. The fifth property is also immediate since we do not create a new cycle when replacing $A$ with $B$.
To see the third property, before replacing $A$ with $B$, we had $A\cup B$ in $\mathcal{C}_{\ell}$, and their exponent was $2^{\ell}$. After the replacement, all cycles in $B\setminus A$ come to have exponent $2\cdot 2^{\ell}=2^{\ell+1}$, the cycles in $A\setminus B$ disappear from $\mathcal{C}_{\ell}$, and those in $A\cap B$ remain unchanged. Note that every cycle remains to have an exponent that is a power of two. Therefore, we have $N_{\ell+1}(\mathcal{C}^{\prime})>N_{\ell+1}(\mathcal{C})$, and $N_{\ell^{\prime}}(\mathcal{C}^{\prime})=N_{\ell^{\prime}}(\mathcal{C})$ for all $\ell^{\prime}\geq\ell+2$.
Observe that due to the second property and the third, the process must terminate. Thus, starting from $\mathcal{C}$, at the termination we have $\mathcal{C}^{\prime}$ that satisfies the first, second, and fourth properties. We know $N(\mathcal{C}^{\prime})=N(\mathcal{C})>4p(\lg L+1)$. Further, we know that cycles of exponent at most $p$ contribute to $N(\mathcal{C}^{\prime})$ by at most $2(\lg L+1)(1+2+4+\ldots+2^{\lfloor\lg p\rfloor})<4p(\lg L+1)$. Thus, there must exist a cycle in $\mathcal{C}^{\prime}$ of exponent greater than $p$. Let $C$ denote the cycle. So, we have shown that
$\Phi(W)=\Phi(T)\Phi(\mathcal{C}^{\prime})\Phi(T^{\prime})=\Phi(T)\Phi(C)^{q}\Phi(\mathcal{C}^{\prime}\setminus C^{q})\Phi(T^{\prime}),$
where $q\geq p+1$. Here $\mathcal{C}^{\prime}\setminus C^{q}$ implies the resulting collection of cycles we obtain after removing $q$ copies of $C$ from $\mathcal{C}^{\prime}$. Now consider walk $W_{q^{\prime}}$ that concatenates $T$, $q^{\prime}$ copies of $C$, $\mathcal{C}^{\prime}\setminus C^{q^{\prime}}$, and $T^{\prime}$. Here, the walk starts with $T$ and ends with $T^{\prime}$, and the cycles can be placed in an arbitrary order. This is because $T$ is a walk from $i$ to $v$, and all the cycles in $\mathcal{C}^{\prime}$ start from $v$ and end at $v$—due to the fifth property—and $T^{\prime}$ is a walk from $v$ to $j$ Further, they are all shorter than $W$ because $W_{q}$ is no longer than $W$ due to the second property. Therefore, $W_{0},W_{1},\ldots W_{q-1}\in\mathcal{W}_{i,j}^{(k)}$. Thus, thanks to $p$-stability, we have $\Phi(W)+A_{i,j}^{(k)}=A_{i,j}^{(k)}$ as desired.
Although we gave the full proof of Theorem 1.5, to convey intuition better, we also give some warm-up analyses in Appendix B by giving a looser bound for the general case and subsequently by considering a special case of $p=1$.
6 Related Work
If the semiring is naturally ordered222 $\bm{S}$ is naturally ordered if the relation $x\preceq_{S}y$ defined as $\exists z:x\oplus z=y$ is a partial order. , then the least fixpoint of a $\text{\sf Datalog}^{\circ}$ program is the least fixed point of $f$ under the same partial order extended to $S^{n}$ componentwise. This is the least fixpoint semantics of a $\text{\sf Datalog}^{\circ}$ program. The naïve evaluation algorithm for evaluating Datalog programs extends naturally to evaluating $\text{\sf Datalog}^{\circ}$ programs: starting from $\bm{x}=0^{n}$, we repeatedly apply $f$ to $\bm{x}$ until a fixpoint is reached $\bm{x}=f(\bm{x})$. The core semiring of a POPS is naturally ordered. Thus, we can find the least fixpoint of a $\text{\sf Datalog}^{\circ}$ program by applying the naïve evaluation algorithm [12].
Computing the least fixpoint solution to a recursive $\text{\sf Datalog}^{\circ}$ program boils down to solving fixpoint equations over semirings. In particular, we are given a multi-valued polynomial function $f:S^{n}\to S^{n}$ over a commutative semiring, and the problem is to compute a (pre-) fixpoint of $f$, i.e. a point $x\in S^{n}$ where $x=f(x)$. As surveyed in in [12], this problem was studied in a very wide range of communities, such as in automata theory [14], program analysis [4, 18], and graph algorithms [3, 17, 16] since the 1970s. (See [19, 9, 15, 8, 23] and references thereof).
When $f=Ax+b$ is linear, as shown in the paper $f^{(k)}(x)=A^{(k-1)}b$ and thus at fixpoint the solution is $A^{(\omega)}b=\lim_{k\to\infty}A^{(k-1)}b$, interpreted as a formal power series over the semiring. If there is a finite $k$ for which $A^{(k)}=A^{(k+1)}$, then it is easy to see that $A^{(\omega)}=A^{(k)}$. The problem of computing $A^{(\omega)}$ is called the algebraic path problem [19], which unifies many problems such as transitive closure [21], shortest paths [5], Kleene’s theorem on finite automata and regular languages [13], and continuous dataflow [4, 11]. If $A$ is a real matrix, then $A^{(\omega)}=I+A+A^{2}+\cdots$ is exactly $(I-A)^{-1}$, if it exists [2, 20, 7].
There are several classes of solutions to the algebraic path problem, which have pros and cons depending on what we can assume about the underlying semiring (whether or not there is a closure operator, idempotency, natural orderability, etc.). We refer the reader to [8, 19] for more detailed discussions.
Appendix A Convergence of Naturally Ordered Semirings
Definition A.1 (Natural Order).
In any (pre-)semiring $\bm{S}$, the relation $x\preceq_{S}y$ defined as $\exists z:x\oplus z=y$, is a preorder, which means that it is reflexive and transitive, but it is not anti-symmetric in general. When $\preceq_{S}$ is anti-symmetric, then it is a partial order, and is called the natural order on $\bm{S}$; in that case we say that $\bm{S}$ is naturally ordered.
For simplicity, we may use $\leq$ in lieu of $\preceq$. We naturally extend the ordering to vectors and matrices: for two vectors $\bm{v},\bm{w}\in\bm{S}^{n}$, we have $\bm{v}\leq\bm{w}$ iff $\bm{v}_{i}\leq\bm{w}_{i}$ for all $i\in[n]$. Similarly, for two matrices $A$ and $B$ (which includes vectors) $A\leq B$ means that componentwise each entry in $A$ is at most the entry in $B$, that is for all $i$ and $j$ it is the case that $A_{i,j}\leq B_{i,j}$.
Here we take $k$ to be the longest chain in the natural order.
Theorem A.2.
Every linear $\text{\sf Datalog}^{\circ}$ program over a $p$-stable naturally-ordered commutative semiring with maximum chain size $k$ converges in $O(kn)$ steps.
Theorem A.3.
There are linear $\text{\sf Datalog}^{\circ}$ programs over a $p$-stable naturally-ordered commutative semiring with maximum chain size $k$ that require in $\Omega(kn)$ steps to converge.
This section considers bounds in terms of the longest chain in the partial order of a naturally ordered semiring. Recall that natural ordering means the following: for two elements $a$ and $b$ $a\leq b$ if and only if there exists a $c$ such that $a+c=b$. Let $L$ be the length of the longest chain in this partial order. We seek to bound the convergence rate in terms of $n$ and $L$.
Lemma A.4.
Consider a naturally ordered semiring where $L$ is the length of the longest chain in the partial order. Let $A$ be an $n\times n$ matrix. Convergence must occur within $nL$ steps.
Proof A.5.
Consider $A^{(k)}x$ as $k$ increases for any fixed $x$. If there is a $k\leq nL$ such that $A^{(k)}x=A^{(k+1)}x$ then convergence has been reached within the desired number of steps. Otherwise when $A^{(k)}x\neq A^{(k+1)}x$ there exists an $i$ such that dimension $i$ in $A^{(k+1)}x$ is strictly greater than dimension $i$ in $A^{(k)}x$. This can only occur $L$ times for each $i$ by definition of the partial order. Knowing that there are at most $n$ dimensions in $A^{(k)}x$, the lemma follows.
Lemma A.6.
There exists a naturally ordered semiring where $L$ is the length of the longest chain in the partial order and a $n$ by $n$ matrix where convergence requires $\Omega(nL)$ steps.
Proof A.7.
Consider the following semiring. The semiring is on the set of integers $0,1,2,\ldots,L$ and a special element $\mathcal{O}$. Here, the additive identity is $\mathcal{O}$ and the multiplicative identity is 0. Consider two elements $a$ and $b$ that are not $\mathcal{O}$. Define the addition and multiplication of $a$ and $b$ to be equal to $\min\{a+b,L\}$. Define $a$ multiplied by $\mathcal{O}$ to be $\mathcal{O}$ for any $a$ and $a$ added to $\mathcal{O}$ to be $a$ for any $a$. Intuitively, addition and multiplication act as standard addition capped at $L$, except for the special $\mathcal{O}$ element.
Consider the following graph corresponding to a $n\times n$ matrix $A$. There is a cycle on $n$ nodes. Order the edges from $1$ to $n$. Each edge is labeled 00 except the edge from $1$ to $2$, which is labeled as $1$. Traversing the cycle $k$ times and multiplying the labels of the edges returns $\min\{k,L\}$. It takes a walk of length $\Omega(nL)$ to reach the element $L$.
Appendix B Warm-up for Proof of Theorem 1.5
In Section 5 we gave the full proof of Theorem 1.5, which gives an upper bound of $O(np\log L)$ on the convergence rate when the underlying semiring has a ground set of size at most $L$.
To convey intuition better of the analysis, we give two warm-up proofs. Our first warm-up is giving a looser bound on the convergence rate. The proof makes use of the fact that a sufficiently long walk must visit the same vertex many times with the same product value. Here, we think of a prefix of the walk as a product of edges on the prefix. This prefex evaluates to an element of the semiring.
Lemma B.1.
Let $A$ be an $n$ by $n$ matrix over a $p$-stable semiring on a ground set $S$ consisting of $L$ elements. Then $A^{(k+1)}=A^{(k)}$, where $k=npL$.
Proof B.2.
Fix $i,j\in[n]$. Consider any $W\in\mathcal{W}_{i,j}^{k+1}$. To show the lemma it suffices to show $\Phi(W)+A_{i,j}^{(k)}=A_{i,j}^{(k)}$. Let $W_{h}$ be the prefix of $W$ of length $h$. Let $v(W_{h})$ denote the ending point of $W_{h}$. Consider all pairs $(\Phi(W_{h}),v(W_{h}))$, $h\in[k+1]$. Since these tuples are subsets of $S\times[n]$ and $|S|=L$, due to the pigeonhole principle, there must exist $H\subseteq[k+1]$ of size $p+1$ such that $(\Phi(W_{h}),v(W_{h}))$ is the same tuple for all $h\in H$. By renaming we can represent the prefixes as $X_{1}$, $X_{1}X_{2}$, $X_{1}X_{2}X_{3}$, $\ldots$, $X_{1}X_{2}X_{3}\ldots X_{p+1}$.
For some $T$ (possibly empty), we have $W=X_{1}X_{2}X_{3}\ldots X_{p+1}T$. By definition $\Phi(X_{1})=\Phi(X_{1}X_{2})=\ldots=\Phi(X_{1}X_{2}\ldots X_{p+1})$.
Thus, $\Phi(W)=\Phi(X_{1}X_{2}X_{3}\ldots X_{p+1}T)$ $=\Phi(X_{1}X_{2}X_{3}\ldots X_{p}T)$ $\ldots=\Phi(X_{1}T)$, This uses the fact that $X_{1}T$, $X_{1}X_{2}T$, $\ldots$, $X_{1}X_{2}X_{3}\ldots X_{p}T$ all are walks from $i$ to $j$ since $X_{1}$, $X_{1}X_{2}$, …, $X_{1}X_{2}\ldots X_{p+1}$ all end where $T$ starts. Further, all walks $X_{1}T$, $X_{1}X_{2}T$, $\ldots$, $X_{1}X_{2}X_{3}\ldots X_{p}T$ are strictly shorter than $W$. This implies that $\Phi(W)$ appears at least $p$ times in $A^{(k)}_{i,j}$. Using Proposition 2.2, we conclude $\Phi(W)+A_{i,j}^{(k)}=A_{i,j}^{(k)}$ as desired.
Next we consider the special case of $p=1$. In this case we give an exponential improvement over what we showed in the previous lemma. The key idea is the following. Previously we identified $p$ disjoint cycles $X_{2},X_{3},\ldots,X_{p+1}$ that share the same starting and ending vertex from a long walk $W$ in $\mathcal{W}^{k+1}_{i,j}$. Then, by removing them sequentially we were able to obtain $p$ copies of the same element that have already appeared; thus adding $W$ (or more precisely $\Phi(W)$) doesn’t change $A^{(k)}_{i,j}$. Now we would like to make the same argument with an exponentially smaller number of cycles. Roughly speaking, we will identify $\Theta(\lg L)$ such cycles and find $2^{\Theta(\lg L)}$ walks by combining subsets of them. That is, the key idea is that we find more walks with the same product from far fewer cycles.
Lemma B.3.
Let $A$ be an $n$ by $n$ matrix both over a $1$-stable semiring $S$ with a ground set consisting of $L$ elements. Then $A^{(k+1)}=A^{(k)}$, where $k=O(n\lg L)$.
Proof B.4.
As before, fix $i,j\in[n]$. Consider any $k\geq\lceil 2\lg L\rceil n$. For any $W\in\mathcal{W}_{i,j}^{k+1}$ we show $\Phi(W)+A_{i,j}^{(k)}=A_{i,j}^{(k)}$. Since there are $n$ vertices, the walk must visit some vertex at least $\lceil 2\lg L\rceil+1$ times. Formally, we can decompose $W$ into
$W=TC_{1}C_{2}\ldots C_{H}T^{\prime}$ (13)
where $T$, $TC_{1}$, $TC_{1}C_{2}$, …, $TC_{1}C_{2}\ldots C_{H}$ all end at the same vertex $v$, and $H=\lceil 2\lg L\rceil$. Note that all the cycles (or closed walks) $C_{1},C_{2},\ldots C_{H}$ start from $v$ and end at the same vertex $v$. It is plausible that some of them are identical.
For a subset $A$ of $[H]$ we let $\hat{\Phi}(A):=\prod_{h\in A}\Phi(C_{h})$. Since there are $2^{|H|}$ subsets of $[H]$ and $2^{H}>L$, there must exist $A,B\subseteq[H]$ such that $A\neq B$ and $\hat{\Phi}(A)=\hat{\Phi}(B)$. Assume wlog that $B\setminus A\neq\emptyset$. Thus we know
$\hat{\Phi}(A\cap B)\hat{\Phi}(A\setminus B)=\hat{\Phi}(A\cap B)\hat{\Phi}(B\setminus A)$ (14)
We can then show,
[…]
$=\Phi(T)\hat{\Phi}(A\cap B)\hat{\Phi}(A\setminus B)\hat{\Phi}(B\setminus A)\hat{\Phi}([H]\setminus(A\cup B))\Phi(T^{\prime})$
[…]
Consider a walk $W^{\prime}$ that starts with $T$, has $C_{h}$ for each $h\in(A\cap B)\cup(B\setminus A)\cup([H]\setminus(A\cup B))=[H]\setminus(A\setminus B)$ and ends with $T^{\prime}$. Similarly, consider a walk $W^{\prime\prime}$ that starts with $T$, has $C_{h}$ for each $h\in(A\cap B)\cup([H]\setminus(A\cup B))$ and ends with $T^{\prime}$. Note that $W$ and $W^{\prime}$ are different since $B\setminus A\neq\emptyset$. Further they are walks from $i$ to $j$ since every cycle $C_{h}$, $h\in[H]$ starts from and ends at the same vertex $v$. Further, both walks are shorter than $W$, and therefore are in $\mathcal{W}^{(k)}_{i,j}$. Since we have $\Phi(W^{\prime})=\Phi(T)\hat{\Phi}(A\cap B)\hat{\Phi}(B\setminus A)\hat{\Phi}([H]\setminus(A\cup B))\Phi(T^{\prime})$ and $\Phi(W^{\prime\prime})=\Phi(T)\hat{\Phi}(A\cap B)\hat{\Phi}([H]\setminus(A\cup B))\Phi(T^{\prime})$. We will show using 1-stability of the semiring that $\Phi(W^{\prime})+\Phi(W^{\prime\prime})+\Phi(W)=\Phi(W^{\prime})+\Phi(W^{\prime\prime})$, implying $\Phi(W)+A_{i,j}^{(k)}=A_{i,j}^{(k)}$ as desired.
Thus, it suffices to show $\Phi(W^{\prime})+\Phi(W^{\prime\prime})+\Phi(W)=\Phi(W^{\prime})+\Phi(W^{\prime\prime})$. To see this:
$\Phi(W^{\prime})+\Phi(W^{\prime\prime})+\Phi(W)$
$=\Phi(T)\hat{\Phi}(A\cap B)\hat{\Phi}(B\setminus A)\hat{\Phi}([H]\setminus(A\cup B))\Phi(T^{\prime})+\Phi(T)\hat{\Phi}(A\cap B)\hat{\Phi}([H]\setminus(A\cup B))\Phi(T^{\prime})$
$\qquad+\Phi(T)\hat{\Phi}(A\cap B)(\hat{\Phi}(B\setminus A))^{2}\hat{\Phi}([H]\setminus(A\cup B))\Phi(T^{\prime})$
$=\Phi(T)\hat{\Phi}(A\cap B)\Phi([H]\setminus(A\cup B))\Phi(T^{\prime})(1+\Phi(B\setminus A)+\Phi(B\setminus A)^{2})$
[…]
$=\Phi(W^{\prime})+\Phi(W^{\prime\prime})$